Using an example, compare and contrast simple interest and compound interest. Please make this answer unique and not like the others that have answered this question.
Answer #1

In simple interest, the interest in each period = principal * periodic interest rate

The interest remains the same in each period, as the principal is unchanged.

However, with compound interest, the interest in each period is added to the principal, and the interest in the next period is calculated on the increased principal. Thus, the interest increases in each successive period as the principal is increased.

An example of simple interest is shown below. The interest rate is 5%, and the starting principal is $1000.

C ДА В Principal at start of 1 Period period 1 $ 1,000.00 2 $ 1,000.00 3 $ 1,000.00 4 $ 1,000.00 5 $ 1,000.00 6 $ 1,000.00 7

Principal at end of period 21 -B2 =B3 AB Principal at start of 1 Period period 1000 3 2 =D2 4 3 =D3 5 4 =D4 65 =D5 =D6 87 =D7

An example of compound interest is shown below. The interest rate is 5%, and the starting principal is $1000.

A B C 1 Principal at Principal at start of end of 14 Period period Interest period 1 $1,000.00 $50.00 $1,050.00 2 $1,050.00 $

14 Period 15 1 16 2 18 4 195 Principal at start of period Interest 1000 =B15*5% =D15 =B16*5% =D16 =B17*5% =D17 =B18*5% =D18 =